OpenStudy (anonymous):

s.one plz help me .! 2cos^2x-2cos2x+4sinxcosx+1 = 0

ganeshie8 (ganeshie8):

\(\cos 2x = \cos^2x - \sin^2 x\)

OpenStudy (anonymous):

i have sin^2x + 2sin2x +1 .... and i stuck

OpenStudy (anonymous):

use 2 cos^2 x = 1-sin 2x after that convert all of them to half angles of tan